FOURIER TRANSFORM INFRARED EMISSION SPECTRA OF ZnH AND ZnH_{2}

A. SHAYESTEH, \underline{D. R. T. APPADOO}, I. GORDON, and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, ON, N2L 3G1, Canada.

We have recorded the vibration-rotation spectrum of gaseous ZnH and ZnH_{2} in emission using a furnace-discharge source. The molecules were generated by flowing a few torr of an Ar/H_{2} gas mixture over a sample of zinc metal in a furnace at 470^{o}C and applying a 3 kV / 333 mA discharge. The observed spectrum contained the fundamental and first hot band of 64ZnH, ^{66}ZnH and ^{68}ZnH in the ^{2}Σ^{+} ground electronic state, the antisymmetric stretching band, 001-000, of ^{64}ZnH_{2}, 66ZnH_{2} and ^{68}ZnH_{2} and the 011-010 hot band of ^{64}ZnH_{2}. The highest signal-to-noise ratio for ZnH lines was about 50. The data analysis is in progress and will be presented.